OpenStudy (nathanjhw):

Consider the function: f(x) =3/4x^4 - x^3 -3x^2 + 6x Determine the interval(s) where f(x) is increasing (if any) and the interval(s) where f(x) is decreasing (if any).

OpenStudy (nathanjhw):

@nincompoop

myininaya (myininaya):

\[f(x)=\frac{3}{4}x^4-x^3-3x^2+6x \\ \text{ First step is to differentiate }\]

myininaya (myininaya):

Then find when f'=0

myininaya (myininaya):

Then we will draw a number line and test intervals.

myininaya (myininaya):

Can you find f'?

OpenStudy (nathanjhw):

f' = 3x^3 -3x^2 -6x +6

myininaya (myininaya):

ok and then we need to find when f'=0

myininaya (myininaya):

it looks like you can factor by grouping if i'm not mistaken

OpenStudy (nathanjhw):

3(x-1)(x^2-2)

myininaya (myininaya):

\[f'(x)=3x^2(x-1)-6(x-1) \\ f'(x)=(x-1)(3x^2-6)\] now set both factors equal to zero and solve to find when f'=0

myininaya (myininaya):

yes that factorization works awesome too

myininaya (myininaya):

I left my 3's either way

OpenStudy (nathanjhw):

x = 1 and x =+- sqrt2

myininaya (myininaya):

 we have 4 intervals to check